OpenStudy (kenttopia2015):

im doing an EOC review an i need help doing a certain problem. What is the Value of this expression when a=64 and b=-5 and i cant type the problem because it has a square root and i dont know hwo to type it on

OpenStudy (kenttopia2015):

\[-2\sqrt[3]{64-5^2}\]

OpenStudy (kenttopia2015):

and ik the answer is 17

OpenStudy (kenttopia2015):

but im not looking for the final solution

OpenStudy (kenttopia2015):

my choices are \[-2\sqrt[3]{89}\] \[-2\sqrt[2]{89}\] \[-2\sqrt[2]{39}\] \[-2\sqrt[3]{39}\]

Nnesha (nnesha):

what's the original expression ?

OpenStudy (kenttopia2015):

the fist expression i typed

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Kenttopia2015 \[-2\sqrt[3]{64-5^2}\] \(\color{blue}{\text{End of Quote}}\) b is -5 so you should put the parentheses \[-2\sqrt[3]{64-(-5)^2}\]

OpenStudy (kenttopia2015):

sorry

OpenStudy (kenttopia2015):

didnt think

Nnesha (nnesha):

is there a plus or negative sign between a and b ???

OpenStudy (kenttopia2015):

i currently dont have a calculator that is able to solve an expression like that

OpenStudy (kenttopia2015):

negative sign

OpenStudy (kenttopia2015):

wait

Nnesha (nnesha):

okay all you need to do is take the the square of 5 and then subtract

OpenStudy (kenttopia2015):

im sorry there is a plus

Nnesha (nnesha):

can you take a screen shot please

OpenStudy (kenttopia2015):

its on an ipad

Nnesha (nnesha):

\[\rm -2\sqrt{a+b^2}\] is this your question ?

OpenStudy (kenttopia2015):

yeah

OpenStudy (kenttopia2015):

wait

OpenStudy (kenttopia2015):

its cubed to the right if the square root

OpenStudy (kenttopia2015):

i mean left

OpenStudy (kenttopia2015):

god why am i so confused about this

Nnesha (nnesha):

ohh yeah sorry

Nnesha (nnesha):

\[\large\rm -2 \sqrt[3]{a+b^2}\] replace a with 64 and `b` with -5

OpenStudy (kenttopia2015):

yeah and you get \[-2\sqrt[3]{64+(-5)}\]

Nnesha (nnesha):

yes don't forget the square yeah \[-2\sqrt[3]{64+(-5)^2}\] take the square (-5)^2= ??

OpenStudy (kenttopia2015):

oh sorry

OpenStudy (kenttopia2015):

25

Nnesha (nnesha):

right \[-2\sqrt[3]{64+25}\] now just add 64+25

OpenStudy (kenttopia2015):

89

OpenStudy (kenttopia2015):

ohhhhhh

OpenStudy (kenttopia2015):

thank you so very much

Nnesha (nnesha):

np

OpenStudy (kenttopia2015):

does the cube stay to the left of the square root?

OpenStudy (kenttopia2015):

nvm

OpenStudy (kenttopia2015):

thanks

Nnesha (nnesha):

after looking at the choices i think we don't have to do anything with cube root